OpenStudy (anonymous):

The value of cotπ/20 cot3π/20 cot5π/20 cot7π/20 cot9π/20 is A) -1 b) 1/2 c) 0 d) 1

OpenStudy (anonymous):

cotπ/20 * cot9π/20 = 1

OpenStudy (anonymous):

cot3π/20 *cot7π/20 = 1

OpenStudy (anonymous):

So....u Need to find only cot5π/20

OpenStudy (anonymous):

cot5π/20 = cot45 = 1

OpenStudy (anonymous):

so 1*1*1 = 1

OpenStudy (anonymous):

@Yahoo! thank you